# PermitQA: A Benchmark for Retrieval Augmented Generation in Wind Siting and Permitting domain

### v1

@@ Abstract
In the rapidly evolving landscape of Natural Language Processing (NLP) and text generation, the emergence of Retrieval Augmented Generation (RAG) presents a promising avenue for improving the quality and reliability of generated text by leveraging information retrieved from user specified database. Benchmarking is essential to evaluate and compare the performance of the different RAG configurations in terms of retriever and generator, providing insights into their effectiveness, scalability, and suitability for the specific domain and applications. In this paper, we present a comprehensive framework to generate a domain relevant RAG benchmark. Our framework is based on automatic question-answer generation with Human (domain experts)-AI (Large Language Model (LLM)) teaming. As a case study, we demonstrate the framework by introducing PermitQA, a first-of-its-kind benchmark on the wind siting and permitting domain which comprises of multiple scientific documents/reports related to environmental impact of wind energy projects. Our framework systematically evaluates RAG performance using diverse metrics and multiple question types with varying complexity level. We also demonstrate the performance of different models on our benchmark.
PermitQA: A Benchmark for Retrieval Augmented Generation in Wind Siting and Permitting domain
Rounak Meyur, Hung Phan, Sridevi Wagle, Jan Strube, Mahantesh Halappanavar, Sameera Horawalavithana Anurag Acharya, Sai Munikoti Pacific Northwest National Laboratory Richland, WA 99354 {rounak.meyur, hung.phan, sridevi.wagle, jan.strube, Mahantesh.Halappanavar, yasanka.horawalavithana, anurag.acharya,sai.munikoti}@pnnl.gov
@@ 1 Introduction
In this work, we present a hybrid workflow to benchmark RAG s, which combines rapid question generation through automated methods, augmented with properly designed human prompts to generate diverse set of questions. Our proposed benchmarking framework is used to generate questions from documents related to wind turbine siting and permitting. These questions serve as a tool to evaluate the performance of RAG-based LLM s, which are designed to answer queries related to these extensive and comprehensive documents. Given the vast amount of information contained in these documents, manually reviewing them is impractical, making RAG-based LLM s essential for generating accurate responses to specific queries. Our benchmarking framework assesses the effectiveness of these models in accurately retrieving and responding to queries, ensuring that they can reliably process and provide relevant information from the documents.
Contributions  The paper introduces a novel benchmark in a specific domain and also proposes a generic framework to evaluate the performance of RAG-based LLM s in responding to different types of questions. This framework is designed to be adaptable across various domains, with a specific focus on wind energy-related documents in this study. The contributions of this research are as follows: (i) We present PermitQA,  [^1] This benchmark will be made publicly available. the first benchmark in the Wind Siting and Permitting domain, (ii) our proposed framework is domain-agnostic, so it can be tailored for any desired niche domain (iii) we introduce a hybrid method to automatically generate various types of questions, producing both objective and subjective responses. The framework also generates questions from different sections of documents to evaluate LLM performance across various sections and question types, and (iv) we utilize existing scoring frameworks like $\mathsf{RAGAS}$ to evaluate model performance, incorporating different LLM s as evaluators for scoring. This approach ensures scalability and quick reproducibility of this approach, while also providing a holistic comparison of LLM performance in terms of responding to questions and assessing or comparing LLM responses with the ground truth answers.
@@ 4 Methodology
Table 2: Performance of the models on the PermitQA benchmark scored using the RAGAS framework across evaluators. The "Prec." and "Rec." mean Context Precision and Context Recall respectively, while "Type" refers to the Question Type. The best performance for each question type per evaluator is highlighted in bold.
@@ 5 Results and Discussion
We evaluate three RAG-based LLM s, namely GPT-4, Gemini, and Claude, on our PermitQA benchmark. The $\mathsf{RAGAS}$ framework is employed for this evaluation, utilizing an evaluator LLM to assess the models’ performance. The assessment includes metrics such as answer correctness, context precision, and context recall, providing a comprehensive understanding of each model’s capabilities in retrieving and generating accurate information from the given context. In our case, we have used GPT-4 and Gemini-1.5Pro as choices for the evaluator LLM s. Figure 4 presents the answer correctness score, while context precision and context recall depicted in Table 2 show the ability of the models to retrieve the context accurately.
@@ 6 Conclusion
In conclusion, this paper presents a versatile framework for evaluating the performance of RAG-based LLM s across various question types and document sections. We showcase this by introducing a hybrid, automated question-generation method that ensures comprehensive coverage of both objective and subjective queries, and implement this for the use case of wind energy related document and present the PermitQA benchmark, which is a first of its kind benchmark in the wind Siting and Permitting space. However, the usefulness of our work goes beyond this niche domain as our approach is domain-agnostic, meaning it can be used for creating benchmark for any domain. Additionally, our use of the RAGAS scoring framework comes with multiple benefits; it allows for a thorough evaluation of model performance, offering a holistic assessment of LLM capabilities, while also having the advantage of being easy for other researchers to adapt this approach for their own work.
@@ 9 Acknowledgment
This research was sponsored by US DOE Wind Energy Technologies (WET) Office under contract XXXXX. This work was done at Pacific Northwest National Laboratory, a multi-program national laboratory operated by Battelle for the U.S. Department of Energy under contract DE-AC05-76RLO1830. This article has been cleared by PNNL for public release as PNNL-SA-xxxxxx.

### v2

# WeQA: A Benchmark for Retrieval Augmented Generation in Wind Energy Domain

@@ Abstract
In the rapidly evolving landscape of Natural Language Processing (NLP) and text generation, the emergence of Retrieval Augmented Generation (RAG) presents a promising avenue for improving the quality and reliability of generated text by leveraging information retrieved from user specified database. Benchmarking is essential to evaluate and compare the performance of the different RAG configurations in terms of retriever and generator, providing insights into their effectiveness, scalability, and suitability for the specific domain and applications. In this paper, we present a comprehensive framework to generate a domain relevant RAG benchmark. Our framework is based on automatic question-answer generation with Human (domain experts)-AI (Large Language Model (LLM)) teaming. As a case study, we demonstrate the framework by introducing WeQA, a first-of-its-kind benchmark on the wind energy domain which comprises of multiple scientific documents/reports related to environmental impact of wind energy projects. Our framework systematically evaluates RAG performance using diverse metrics and multiple question types with varying complexity level. We also demonstrate the performance of different models on our benchmark.
WeQA: A Benchmark for Retrieval Augmented Generation in Wind Energy Domain
Rounak Meyur, Hung Phan, Sridevi Wagle, Jan Strube, Mahantesh Halappanavar, Sameera Horawalavithana, Anurag Acharya, Sai Munikoti Pacific Northwest National Laboratory Richland, WA 99354 {rounak.meyur, hung.phan, sridevi.wagle, jan.strube, mahantesh.halappanavar, yasanka.horawalavithana, anurag.acharya, sai.munikoti}@pnnl.gov
@@ 1 Introduction
In this work, we present a hybrid workflow to benchmark RAG s, which combines rapid question generation through automated methods, augmented with properly designed human prompts to generate diverse set of questions. Our proposed benchmarking framework is used to generate questions from documents related to wind energy domain. These questions serve as a tool to evaluate the performance of RAG-based LLM s, which are designed to answer queries related to these extensive and comprehensive documents. Given the vast amount of information contained in these documents, manually reviewing them is impractical, making RAG-based LLM s essential for generating accurate responses to specific queries. Our benchmarking framework assesses the effectiveness of these models in accurately retrieving and responding to queries, ensuring that they can reliably process and provide relevant information from the documents.
Contributions  The paper introduces a novel benchmark in a specific domain and also proposes a generic framework to evaluate the performance of RAG-based LLM s in responding to different types of questions. This framework is designed to be adaptable across various domains, with a specific focus on wind energy-related documents in this study. The contributions of this research are as follows: (i) We present WeQA,  [^1] This benchmark will be made publicly available. the first benchmark in the Wind Energy domain, (ii) our proposed framework is domain-agnostic, so it can be tailored for any desired niche domain (iii) we introduce a hybrid method to automatically generate various types of questions, producing both objective and subjective responses. The framework also generates questions from different sections of documents to evaluate LLM performance across various sections and question types, and (iv) we utilize existing scoring frameworks like $\mathsf{RAGAS}$ to evaluate model performance, incorporating different LLM s as evaluators for scoring. This approach ensures scalability and quick reproducibility of this approach, while also providing a holistic comparison of LLM performance in terms of responding to questions and assessing or comparing LLM responses with the ground truth answers.
@@ 4 Methodology
Table 2: Performance of the models on the WeQA benchmark scored using the RAGAS framework across evaluators. The "Prec." and "Rec." mean Context Precision and Context Recall respectively, while "Type" refers to the Question Type. The best performance for each question type per evaluator is highlighted in bold.
@@ 5 Results and Discussion
We evaluate three RAG-based LLM s, namely GPT-4, Gemini, and Claude, on our WeQA benchmark. The $\mathsf{RAGAS}$ framework is employed for this evaluation, utilizing an evaluator LLM to assess the models’ performance. The assessment includes metrics such as answer correctness, context precision, and context recall, providing a comprehensive understanding of each model’s capabilities in retrieving and generating accurate information from the given context. In our case, we have used GPT-4 and Gemini-1.5Pro as choices for the evaluator LLM s. Figure 4 presents the answer correctness score, while context precision and context recall depicted in Table 2 show the ability of the models to retrieve the context accurately.
@@ 6 Conclusion
In conclusion, this paper presents a versatile framework for evaluating the performance of RAG-based LLM s across various question types and document sections. We showcase this by introducing a hybrid, automated question-generation method that ensures comprehensive coverage of both objective and subjective queries, and implement this for the use case of wind energy related document and present the WeQA benchmark, which is a first of its kind benchmark in wind energy domain. However, the usefulness of our work goes beyond this niche domain as our approach is domain-agnostic, meaning it can be used for creating benchmark for any domain. Additionally, our use of the RAGAS scoring framework comes with multiple benefits; it allows for a thorough evaluation of model performance, offering a holistic assessment of LLM capabilities, while also having the advantage of being easy for other researchers to adapt this approach for their own work.
@@ 9 Acknowledgment
This research was sponsored by US DOE Wind Energy Technologies Office (WETO) and was done at Pacific Northwest National Laboratory, a multi-program national laboratory operated by Battelle for the U.S. Department of Energy under contract DE-AC05-76RLO1830. This article has been cleared by PNNL for public release as PNNL-SA-203893.